OpenStudy (ffasinger):

How would you find the fraction of a circle that an arc covers? A. Divide the circumference by 360. B. Divide the measure of the arc by 180. C. Divide the diameter by 360 D. Divide the measure of the arc by 360.

OpenStudy (raffle_snaffle):

well we know it's not?

OpenStudy (ffasinger):

I know it is not B

OpenStudy (misssmartiez):

Take a look at it.

OpenStudy (ffasinger):

I just did.

OpenStudy (misssmartiez):

Do you know now?

OpenStudy (misssmartiez):

Or more clarification?

OpenStudy (ffasinger):

A I think

OpenStudy (misssmartiez):

What do you think mathmale? I'd say A also, http://www.nelson.com/school/elementary/mathK8/quebec/0176237879/documents/NM8SB_5A.pdf Since I don't see an answer saying anything about the 'central angle'.

OpenStudy (mathmale):

Please, could we focus on understanding instead of just answers? answers.com will not be available to you during exams. If an arc length of less than one whole circumference is given to you, write the RATIO \[\frac{ given~arc~length }{ 2\pi r }=\frac{ r \theta }{ 2 \pi r }.\]

OpenStudy (ffasinger):

hmmm

OpenStudy (mathmale):

Notice that you can reduce the last result by cancelling out the ' r ' This fraction is the answer to your question, so long as you express your angle (theta) in radians (not in degrees).

OpenStudy (ffasinger):

ok so it is A

OpenStudy (rockycatsrule):

that seems correct to me

OpenStudy (mathmale):

Can you convince me that "A" is the same thing as I've suggested (using that formula)?

OpenStudy (misssmartiez):

I'd say so. The others don't touch on what we went over. Basically the only thing that makes sense. However. do you know WHY it is A?

OpenStudy (mathmale):

Suppose that I tell you that the length of the arc is 5. What fraction of the whole circular area is represented by that arc length and the corresponding, pizza-slice-shaped area?

OpenStudy (mathmale):

Hint: It's more than half the circle but less than the full circle.

OpenStudy (mathmale):

Write your answer as a proper fraction.

OpenStudy (rockycatsrule):

i dont want to give a direct answer

OpenStudy (rockycatsrule):

cuz that would be agenst contract

OpenStudy (mathmale):

Any explanation you'd like to provide, as an alternative to others' answer choices, would be welcome. You don't need to give the answer, but you could certainly explain how you or FFASinger could go about finding it.

OpenStudy (rockycatsrule):

well i could be wrong but everyone makes misstakes

OpenStudy (rockycatsrule):

but u said it would be half so basiced on that formula i got 180

OpenStudy (phi):

one way to figure it out. look at a simple case